OpenStudy (anonymous):

find the coordinates of the midpoint whose endpoints are A (-5,3) and B (3,3)

random231 (random231):

x= (-5+3)/2=-1 y= (3+3)/2=3

OpenStudy (kc_kennylau):

The midpoint of A(x1,y1) and B(x2,y2): \[x=\frac{x_1+x_2}2\]\[y=\frac{y_1+y_2}2\]